find the exact value of (pi/6) using the 6 trig functions
Answer by lwsshak3(11628)

find the exact value of (pi/6) using the 6 trig functions
**
π/6 is in quadrant I where all the trig functions are positive. You will be working with a 30-60 right triangle:
sinx=1/2
cosx=√3/2
tanx=√3/3
cscx=2
secx=2√3/3
cotx=√3